A Complete Bibliography of The X Resource: A Practical Journal of the X Window System

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

20 October 2023 Version 1.28

Abstract

This bibliography records the contents of the quarterly journal The X Resource: A Practical Journal of the X Window System, ISSN 1058-5591 and ISBN 0-937175-79-X.

Correspondence can be addressed to:

The X Resource O’Reilly and Associates, Inc. 103 Morris Street Sebastopol, CA 95472 USA

E-mail:  adrian@ora.com and
uunet!ora!adrian.

Title word cross-reference

10th [Fer95a].

2D [Léo95].

3-D [Mor93].

5 [Sch91, WJ91]. 5.1P [WFB^{+}92, Wom92].

6.0 [Ste93, Ste94].

8th [Ano93a, Ano93b].

’95 [Nye95b]. 9th [Ano94a].

Access [ELMW95, MST95]. Accessible [WNTV93]. Administration [Hes94].

Administrator [BM94, Hes94, MO92, Nye91, O’N92, Pea92]. After [Gil92]. Again [Phe94]. Against [Kar92, Flo94, RM94]. Agents [McG91]. Analysis [MC93]. Animation [Gas93]. Annual [Ano94a, Ano93a, Ano93b, Fer95a]. API [Eli95, Mor93]. Application [Bue92, Cou91, Fla92, Jon92, Nye93, PC92, Van93, Wer94, Szc92]. Application-Level [Wer94]. Applications [AM95, Ais93, Ano93g, Blo92, CG93, FR94, New95a]. Approach [DR94, Nat94]. Architectural [Eva95]. Architecture [Fla93a, Nel95]. Architectures [Jon93, LE95]. Archive [Ano93k]. Audio [BHM^{+}92, FR94, NN93]. Authentication [Yu94]. Automated [Kep92]. Automatic [Ran95, RC94a]. Auxiliary [Fla92]. Available [Ano92b].

Back [Ano93c]. Bandwidth [FK93]. bars [Flo94, RM94]. Based [Beu95, KW93, Pac93, Pac94, Pri94, RBKL92, Coh93, Hes94, Nat94, SV93]. Best [Alb91, Alb92a, Alb94a, Alb94b, Alb95a, Alb95b, Alb92b, Alb93a, Alb93b, Alb93c]. Bet [Alb94c, Mal92]. Better [Dav92b]. Beyond [MST95, Mar94b]. Bibliography [Eps92]. Binding [Ste92]. Blackboard [AM95]. Born [Phe94]. Bringing [Dei94]. Browser [Van93]. Buddy [Sch93a]. Budget [Pac95]. Buffer [Ell95a, MMG94]. Buffering [Gas93]. Building [CG93, Eps91, LP93, Nye91]. built [BSK95]. Bus [Beu95].

C [Bäc92, Eli95, MAS94, Men93, RB93, Ste92]. Call [Ano94a]. Can [Mal92, Sch93a]. Capabilities [DM92, Mey94]. Cases [Bil94]. CDE [KBSW95, Nye95b]. CDE/Motif [KBSW95]. Center [Har92]. Centered [Bil94]. change [Flo94, RM94]. Changes [Mey94, Sch94b]. Class [Hum93]. Classes [Bäc92]. Client [Fah94, Mar94b, Mor94a, RM94, SB94, Sri95, NN93]. Client/Server [NN93]. Clients [Bal93, NP94, WJ91, Cha95]. Clipboard [RC94b]. CMAP [Ano92a]. Co [RBKL92]. Co-Existent [RBKL92]. Code [Ano94c, Ano94d, Ano94e, Ano92f, Ano93h, Ano93k, Ano93i, Ell95b, Fer95e, Fer95f, Gil94, Nye95c]. Collisions [Tab93]. Color [Ber95, Mai94, TT91]. Colormap [CCR91, Mai94]. Combining [Dyk94]. Command [Mal92]. Commercially [Ano92b]. Common [Eva95]. Communication [Mar94b, RM94, BSK95]. Communications [Sha93]. Comparison [Ran94]. compatibility [Kil95]. Components [Pun95]. Computer [Dyk94]. Computing [KW93]. Conference [Ano94a, Ano93a, Ano93b, Nye95b, Fer95a]. Configuration [Hes94]. Connections [Yu94]. Consortium [Ano93j, Flo94, Gil94, GCGW94, MAS94, Mor94a, Mor94b, NH94, RM94, Sch94a, Sch93a, Sch94c, SB94, Wex94a, Mor93, Sch91, Sch92, Sch94b, Sch95a]. Constraint [Pac93]. Constructing [Bue92, Men93]. Construction [Sta95]. Context [Sch93b]. Contrib [Dav93b]. Contributed [Dav93b]. Control [Mal92, NP94]. Controlling [Dyk92]. Conventions [Flo94, RM94]. Conversion [Ran95]. Cooperation [MST95]. CORBA [Pri94]. CORBA-Based [Pri94]. Cornell [Dyk91]. Creation [SN92]. Cross [Kep92]. Cross-platform [Kep92]. Current [NN93, Sch92]. Custom [Gou94]. Cut [Ben94].

D [Mor93]. D11 [Kil95]. Dance [FR94]. Dash [Van93]. Data [Coh93, Fla93a]. Debugging [Nye93]. Default [CCR91]. Defect [Nat94]. Delivery [Dyk92]. Demystifying [Dav92a]. Dependent [CG93]. Describing [Coh93]. Description [Flo94]. Design [Bil94]. Designer [Ano94a].

Designing [Bäc92, McG92, Pac94]. Desktop [Eva95, McG91, Ran94]. Determinism [Che93]. Developers [Nye95b]. Developing [Bre94, WJ91]. Development [Ran94, Sri95, Szc92]. Directory [Ano92b]. Disabilities [Pac92, WNTV93]. Disk [Nye91]. Display [Fla93a, O’N92]. Displays [Mal92]. Distributed [Bal93, LP93, Mai94, Ros93]. Distribution [Gil92]. doin [Ste94]. domain [Con92]. donated [DM92]. Don’t [Bar92]. Double [Ell95a]. Draft [Flo94, GCGW94, MAS94, Mor94a, Mor94b, NH94, RM94, SB94, Ste95, Wex94a, WFB^{+}92, Zim95a, Zim95b]. Drag [BDC92, DR94, RC94b]. Draw [Bar92]. Drawing [Fla93b]. Drop [BDC92, DR94, RC94b].

Easy [Gou94]. Edit [BSK95]. Editor [Ano92c, Ano93d, Ano93e, Ano93f, Ano94b, Fer94a, Fer94b, Fer95b, Fer95c, Fer95d, Nye94, Nye95a, Pet91a]. editres [Pet91a, Pet91b, Pet91c]. Effect [TMM95]. Effortlessly [Gas93]. Efforts [NN93]. Elegant [Gem95]. Embedded [Sta95]. Embedding [New95a, Pri94]. Enable [Mey94]. Encoding [WFB^{+}92, Wom92]. Engine [Men93]. Engineering [Cou91, KBSW95]. Enhancements [DM92]. Environment [BM94, Ben94, Ber95, BHM^{+}92, Kar92, Pin94, Ros93, YYPC95, Eva95]. Environments [Dyk94, Ran94]. Ergonomics [Bil94]. Event [Dyk92, Pin94, Raw92]. Event-Free [Pin94]. Exchange [Mor94a, SB94]. Existent [RBKL92]. Experience [Fah94]. Experiences [Che93]. Experiment [Pac94]. Exploring [Nie95]. Extending [Che93, Pri94, Sch95b, Zim95a, Zim94]. Extension [Asm95, Bre94, Ell95a, Fer92c, Fer92a, Fer92b, For94, Gas93, GCGW94, Hum93, Kar92, Kil92, Pac95, TMM95, Zim95b, Zim95c, Zim95d].

Facility [Raw92, Wer94]. February

[Zim95a, Zim95b]. Fidget [Bar92]. Figures [Fla93b]. File [Nis93]. Files [BM94, Fla92]. Finding [Fla92]. Fixed [PZL94]. Fixed-Pitch [PZL94]. Focus [Dyk92]. FocusMaps [Cha95]. Font [DM92, Flo94, MO92, Mey94, Pea92, PZL94, RG93]. Form [Ber92]. formal [Nat94]. Format [Le 94, PZL94]. Formats [Coh93, MO92]. Foundation [GT91]. Frame [MMG94]. Framework [Léo95, Lin92a]. Free [Pin94, GT91]. Freely [Ano92b]. Fresco [LP93, LTC94]. Fully [Dai93]. Functional [Dai93]. Future [Fah94, Sha94]. fvwm [Sch95b, Que94].

Generation [Ano93j]. Generator [Bue92]. Geometry [Dar92, Hum93]. Get [Ste93]. Goes [MMG94]. GoldenGate [Ran95]. Grabs [Dyk92]. Graphical [KBSW95, Pet91a]. Graphics [Léo95, Mor93, Pin94]. Group [Ano91b]. Groups [Ano92g]. GUI [Ais93, Kep92, Ran94, Sta95, Szc92]. GUIs [Bue92].

Halliburton [Mai94]. Handler [RG93]. HBF [PZL94]. Hdial [Ano92d]. Help [Dav93a, Ell95b, Gem95, Hem93, Pac92, Sch93b, SV93, Van93]. Hierarchical [Pac93]. High [Kar92, Kil95]. high-performance [Kil95]. Hollywood [MMG94]. HP [DM92]. HP-donated [DM92]. Hush [Eli95]. Hypergraphics [Ous93]. Hypertext [Hem93, Ous93].

I/O [RBKL92]. ICE

[BSK95, Beu95, Mor94a, SB94]. ICE-Based [Beu95]. Imaging [Cou91]. Imake [Dav92a]. Imakefile [Dub93]. Impaired [ER93]. Implement [New92]. Implementation [Dai93]. Implementing [BDC92, Ben94, Lin92b, Sch93b]. Improve [Mey94]. Improved [Hum93]. Improvements [BD92, BK93]. Improving [Bil94, PC92]. Inc. [Sch94c]. Increase [Mey94]. Independence [Lin92b]. Information [Fre95]. Input [Fer92a, Fer92b, NS94, NH94]. Insights [Cou91]. instead [Ste94]. Instrumentation [Kil92]. Integrating [Ber95, Kar92]. Intelligent [Pun95]. Inter [Mar94b, Mor94a, RM94, SB94]. Inter-Client [Mar94b, Mor94a, RM94, SB94]. Interaction [Dyk94]. Interactive [Cou91, Ran94]. Interchange [Coh93]. Interface [KBSW95, MAS94, Pun95, RMR93]. Interfaces [ER93, LP93]. Internationalized [WJ91]. Interoperability [Har92]. Interprocess [Sha93]. Interviews [BS93]. Intrinsics [MAS94, RB93]. Introduction [Gas92, Mar94a, Nie95, TT91]. Issues [Ano93c, Eps91].

Java [Nie95].

k-Edit [BSK95]. Keeping [BM94]. Kerberos [Yu94]. Keyboard [Dyk92, For94]. Knvas [Léo95]. Koalatalk [Beu95]. Korean [YYPC95]. Kornshell [Ran94].

Language [MAS94, Ste92, WJ91]. Large [Ros93]. Layer [Wei92]. Layered [Dai93, Fre95]. Layout [Gou94, Pac93]. LBX [FK93, Pac94]. Letters [Ano93g]. Level [Wer94]. Libraries [SWC92]. Library [Le 94, Men93, Mor94a, Mor94b, SN92, Wer94, Zim95b, BSK95]. Licensing [Mey94]. Life [Mal92]. Light [Beu95]. Light-Weight [Beu95]. Limited [Nye91]. Loadable [LE95]. Logical [Flo94]. Looker [Fre95]. Looker-Upper [Fre95]. Low [FK93].

Macro [Raw92]. Made [Gou94]. Make [FR94]. Making [Dav92b, WNTV93]. Man [Phe94]. Management [Ber95, Dar92, Hum93, Lin92a, Mai94, Mor94b, TT91, Wex94b, Wex94a]. Manager [Gou94, O’N92, Que94]. Managing [Pea92, Ros93]. Manipulation [CCR91]. manpage [Pet91b]. Manual [Pac92, RM94]. Measuring [Gen93]. Mechanism [MST95]. MediaWarehouse [Ran95]. Melding [RB93]. Memory [Bal93]. Menus [RC94a]. Mercator [RMR93]. Message [Beu95, Wer94]. Message-Reporting [Wer94]. MetaCard [Ran94]. Method [Gen93, NS94, NH94, Gen93]. Methodologies [Dyk94]. Methodology [Sri95]. Migrating [Jac92]. minute [Gou94]. Mixed [BM94]. Mobile [KW93, Ric95]. Mobility [BH94]. Model [Ber92, RC94b]. Module [Sch95b]. Monitoring [Gen93]. MOTIF [Bue92, Bäc92, Ber92, DR94, Dar94, Gem95, Gou94, Gou93, KBSW95, Nye95b, RB93, RC94a, RC94b, Sch93b]. Movement [SKD94]. MS [Par91]. MS-Windows [Par91]. MT [SWC92]. MT-Safe [SWC92]. Multi [Gas93, Gil93, Jon92, KHLS94, Smi92]. Multi-Buffering [Gas93]. Multi-rendering [KHLS94]. Multi-Threaded [Gil93, Smi92]. Multi-user [Jon92]. Multicasting [BH94]. Multicomputers [Bal93]. Multidisplay [Jon93]. Multimedia [RBKL92]. Multiple [Jon94]. Multiplexing [BH94]. Multithreaded [Sri95]. MuX [RBKL92]. My [Ell95b].

NASA [Szc92]. Near [Ais93]. Near-Real-Time [Ais93]. Nesting [Pac95]. Netnews [Alb94a, Alb94b, Alb94c, Alb95a, Alb95b, Alb93a, Alb93b, Alb93c]. Network [Alb91, Alb92a, Alb92b, FR94]. News [Alb91, Alb92a, Alb92b, Ano91b]. Next [Ano93j]. NFB [Wei92]. Nonrectangular [Fer92c]. Nonvisual [RMR93]. Notes [Par91].

O [RBKL92]. Object [DR94, Raw92]. Object-Oriented [DR94]. Objects [Bil94]. Obtaining [Ano94c, Ano94d, Ano94e, Ano92f, Ano93h, Ano93i, Fer95e, Fer95f, Nye95c]. One [Gou94]. One-minute [Gou94]. Online [Gem95, Hem93, SV93]. OpenDoc [Nel95]. OpenGL [KHLS94]. Optimizing [PZL94]. optional [Kil95]. O’Reilly [McG91]. Oriented [DR94]. OSF [Bäc92, RB93]. OSF/Motif [Bäc92, RB93]. Other [Dav93b]. Overlays [New92]. Overview [Eva95, For94, NS94].

Package [Gen93, RG93]. Page [Dan95]. Pages [Fer92b]. PanoramiX [Asm95]. Papers [Ano94a]. Paradigm [Kep92]. Past [Fah94]. Paste [Ben94]. Pen [KW93]. Pen-Based [KW93]. Pens [KBSW95]. People [Pac92, WNTV93]. Performance [BD92, Kar92, Laz95, Mey94, PC92, Kil95]. PEX [Har92, KHLS94, Sch95c, Ste93, Ste94, Ste95, WFB^{+}92, Wom92]. PEXlib [Gas92, Mor93, Sch95c, Ste92]. PhotoKit [HS95]. Pitch [PZL94]. platform [Kep92]. PLUS [Szc92]. Pop [RC94a]. Pop-up [RC94a]. Portability [Par91]. Portable [Gil94]. Porting [Wei92]. POSC [Mai94]. POSC/Halliburton [Mai94]. PostScript [Dan95, Rav92]. Prescient [McG91]. Present [Fah94]. Previewing [Dan95]. Print [Dei94]. Program [Str93]. Programmers [Pet91a]. Programming [Ano93g, BK93, Blo92, New95b]. Programs [Bue92, Gem95, Sha93]. Project [RMR93]. Projects [Sch92]. Proposal [SB94, SV93]. Protocol [BH94, Bre94, ELMW95, MC93, NS94, NH94, Ste95, SV93, Wex94b, Wex94a, WFB^{+}92, Wom92, Kil95, Pet91c]. protocol-optional [Kil95]. Prototyping [SWC92]. Provide [Che93]. Pseudo [Men93, Pac95]. Pseudo-Root [Pac95]. Pseudo-servers [Men93]. Pseudoserver [SKD94]. Pseudoservers [Kem94]. Public [Con92, Flo94, GCGW94, MST95, MAS94, Mor94a, Mor94b, NH94, RM94, SB94, Ste95, Wex94a, WFB^{+}92, Zim95a, Zim95b]. Public-domain [Con92]. Publish [Ano91a, Ano92e].

R6 [Mey94, Sch94a]. Radar [McG91]. Rationale [For94]. Re [KBSW95]. Re-engineering [KBSW95]. Real [Ais93]. Record [Zim95b]. Recording [Zim95a, Zim94]. Redisplay [LTC94]. Reference [Fer92b]. Release [BD92, Gil92, MAS94, WJ91, Sch91]. Remote [ELMW95]. Rendering [Kar92, Mey94, KHLS94]. Report [Sch95a, Sch94c]. Reporting [Wer94]. Resolution [Lin92b]. Resolving [Tab93]. Resource [Pet91a, Tab93, Ano93c, Ano93k, Ano91a, Ano92e]. Resources [BK93, Dav92b, Mai94, O’N93]. Reusable [Bäc92, Pun95]. Review [Flo94, GCGW94, MAS94, Mor94a, Mor94b, NH94, RM94, SB94, Ste95, Wex94a, WFB^{+}92, Zim95a, Zim95b]. Root [Pac95]. RPC [Ano93g, Blo92, Sch93a]. Ruler [Ber92]. Runtime [ER93].

Safe [SWC92]. Scalable [BH94, Hes94].

Science [MST95]. Screen

[Jon94, Lin92a, O’N93]. Screen-specific [O’N93]. Scripting [Sta95]. Seconds [Dan95]. Security [Eps92]. Selection [RC94b]. semantics [Kil95]. Server [Dei94, DM92, Fah94, Jon94, Kil92, KHLS94, LE95, Rav92, RBKL92, Smi92, Wei92, NN93]. Servers [Pea92, Men93]. Service [Hes94]. Session [Mor94b, Wex94b, Wex94a]. Sessions [Ric95]. Set [Dyk91]. Shape [Fer92c]. Shared [Mai94]. Sharing [BH94].

Shell [New95b]. Simple [BM94]. Simplifying [Sta95]. Simulation [Raw92]. Sing [FR94]. Smart [MMG94]. Smooth [Gas93]. Software [Dav93b, Gil92, Jon92, Jon93, MST95]. Source [Ano94c, Ano94d, Ano94e, Ano92f, Ano93h, Ano93i, Fer95e, Fer95f, Nye95c]. Space [MST95, Nye91]. Spare [Sch93a]. Specific [Raw92, O’N93]. Specification [Ste92, Ste95, Wom92, Zim95b]. Spoken [NP94]. Spy [Ell95b]. Standard [Flo94, GCGW94, MAS94, Mor94a, Mor94b, NH94, Pac94, RM94, SB94, Wex94a]. Start [BM94]. Start-up [BM94]. Statistics [Gen93]. Status [Sch95a, Sch94c]. Structured [Pin94]. Support [ER93, PZL94, Pri94]. Supporting [KW93]. Survey [Jon93, Kem94]. Synchronization [GCGW94, Zim95c]. Synlib [Cha95]. System [Ber95, FR94, Gem95, Hem93, Mai94, MAS94, Mor93, Nis93, RMR93, Van93, Kil95, ELMW95, Laz95, Lin92b, Mal92, MC93, Pac95, TT91, WNTV93, YYPC95]. Systems [Dav93a, Eps92, Eps91].

Table [Har91]. TAE [Szc92]. Tale [BS93]. Talks [Ano94a]. Taming [O’N92]. Task [Lin92a]. TaX [CG93]. Tcl [Eli95, Mar94a, Ran94]. Tcl/Tk [Eli95, Ran94]. Technical [Ano94a, Ano93a, Ano93b, Fer95a, TT91]. Technology [Mey94]. Telephone [Dan95]. Teleporting [Ric95]. Testing [Cha95, Dar92, Kep92, Zim95d, Nat94]. There [Ste93, Ell95b]. Threaded [Gil93, Smi92]. Three [Dav93a]. Tickling [Sch95b]. Time [Ais93, CG93, RBKL92]. Time-Based [RBKL92]. tips [Dav91]. Tk [Eli95, Mar94a, Ous93, Ran94, Sch95b]. TkMan [Phe94]. Tool [CG93, Szc92]. Toolkit [Fla92, HS95, MAS94, Str93, Man93]. Toolkits [BS93, Nat94]. Tools [Gen93]. top [Lin92b]. Trace [MC93]. Tracing [Kil92]. Traitifying [Dar94]. Transfer [RC94b]. Translation [ER93]. transport [Kil95]. transport-optional [Kil95]. Tree [Dav93b]. Trestle [Man93]. tricks [Dav91]. Troubleshooting [Dub93]. True [Jon94]. Trusted [Eps91]. TTY [Bue92]. Tutorial [Fer92a, Fla94, Le 94, Mar94a, Nie95, Zim95d]. Tutorials [Ano94a]. Two [BS93].

Uniform [RC94b]. Unifying [RC94b]. University [Dyk91]. UNIX [Hes94, Ber95]. UNIXhelp [Dav93a]. Update [FK93]. Upper [Fre95]. Usability [Bil94, McG92]. Usage [Gen93]. Use [Bil94, Dav92b]. User [Ano91b, Ano92g, Bil94, Dav91, Dav92a, Dav92b, Dav93b, Fla93b, Fre95, KBSW95, LP93, New95b, O’N93, Phe94, Pun95, Que94, Jon92]. User-Contributed [Dav93b]. Users [ER93, Pet91a]. uses [WJ91]. Using [Cha95, DM92, Jon92, Pac92, Mal92]. Utilities [MO92].

Version [Flo94, MAS94, RM94, Ste95, WFB^{+}92, Wom92, Zim95a, Zim95b]. versus [BS93]. Virtual [Dyk94, Lin92a]. Visual [TMM95]. Visualizing [Laz95, LR92]. Visually [ER93]. Visually-Impaired [ER93].

Washington [Fre95]. Watcher [NP94]. Way [Str93]. Wcl [Har91, SN92]. Web [Dar95]. Weight [Beu95]. Well [FR94]. We’re [Ste94]. Widget [Ano92a, Ano92d, Bäc92, Ber92, Con92, Dar92, Dyk91, Fla94, Fla93a, GT91, Har91, Hum93, Léo95, Pac93, SN92, Gen93]. Widgets [Ano92b, Bar92, Gou94, Gou93, Jac92, Nat94]. Will [Ste93]. WILLOW [Fre95]. Window [BH94, ELMW95, Eps91, Fer92c, Laz95, Lin92b, Mal92, MAS94, Mor93, NP94, Pac95, Que94, RMR93, SKD94, YYPC95, Kil95, MC93, WNTV93]. Windowing [Eps92]. Windows [Dai93, Fre95, Par91]. Words [NP94]. Working [Kar92]. Workstation [BHM^{+}92]. Writing [Gil94, Gou93]. WYSIWYG [Dei94].

X [Ano91a, Ano92e, Fer95a, AM95, Ais93, Ano94a, Ano92g, Ano93c, Ano93g, Ano93j, Ano93k, Ano93a, Ano93b, Bal93, BM94, Ben94, BHM^{+}92, Bil94, Blo92, BH94, Bre94, Cha95, Che93, Coh93, Dar95, Dav91, Dav92a, Dav92b, Dei94, ELMW95, ER93, Eps91, Fla92, Fla93b, Flo94, For94, Fre95, FK93, Gem95, Gil94, Gil92, GCGW94, Hes94, Jon93, Jon94, Kar92, KW93, KHLS94, Laz95, LR92, Lin92b, Mai94, Mal92, MST95, MC93, MAS94, MO92, Men93, Mor93, Mor94a, Mor94b, NH94, NP94, New92, New95a, New95b, Nis93, Nye91, Nye93, O’N92, O’N93, Pac95, Par91, Pea92, PC92, Phe94, Pin94, Que94, Rav92, Raw92, RBKL92, RMR93, Ric95, RM94, Ros93, Sch91, Sch92, Sch94a, Sch94b, Sch95a]. X [Sch93a, Sch94c, SB94, Sha94, Smi92, SKD94, Sri95, Str93, SV93, TT91, TMM95, WNTV93, Wei92, Wex94a, WJ91, WFB^{+}92, YYPC95, Yu94, Zim95a, Zim95c, Zim94]. X-based [Coh93, Hes94, SV93]. X-Clients [Cha95]. X.deskhelp [Dav93a]. X/Motif [Gem95]. X11 [BDC92, Ber95, Dyk94, Fer92a, Fer92b, Kil95, WJ91, Zim95d]. X11R5 [Nye91]. X11R6 [Mar94b, Mey94]. xbb [AM95]. XcRichText [Con92]. Xcu [Dyk91]. xdm [BM94]. xdm/xinit [BM94]. xfig [Fla93b]. XGEN [Bue92]. xhelp [Dav93a]. Xhyper [Hem93]. XIE [Fah94, HS95]. xinit [BM94]. XKB [For94]. Xlib [Gil93]. Xm [Str93]. Xmc [BH94]. XMOVE [SKD94]. Xmsg [Wer94]. XmtLayout [Fla94]. XPM [Le 94]. xrdb [O’N93]. XScript [New95b]. XSMP [Wex94b]. Xt [BS93, BD92, Jon92, Nat94, Pri94, Sha93, SWC92, Tab93]. Xt-based [Nat94]. xtent [BK93]. xterm [Dav91]. XTrap [Pac92]. Xvan [Jon94]. XVE [TMM95]. XView [SWC92]. Xy [BH94].

Zero [Nat94]. Zero-Defect [Nat94].

References

Aisic:XR-5-1-183

[Ais93] Ilan Aisic. GUI for near-real-time applications in X. The X Resource, 5(1):183–191, January 1993. CODEN XRESEA. ISBN 1-56592-020-1. ISSN 1058-5591.

Albert:XR-0-1-17

[Alb91] Marc Albert. Best of network news. The X Resource, 0(1):17– 23, October 1991. CODEN XRE-SEA. ISBN 0-937175-79-X. ISSN 1058-5591.

Albert:XR-1-2-49

[Alb92a] Marc Albert. Best of network news. The X Resource, 2(1):49– 53, April 1992. CODEN XRE-SEA. ISBN 0-937175-97-8. ISSN 1058-5591.

Albert:XR-3-1-47

[Alb92b] Marc Albert. Best of network news. The X Resource, 3(1):47– 54, June 1992. CODEN XRE-SEA. ISBN 0-937175-98-6. ISSN 1058-5591.

Albert:XR-6-1-45

[Alb93a] Marc Albert. Best of netnews. The X Resource, 6(1):45– 48, March 1993. CODEN XRE-SEA. ISBN 1-56592-021-X. ISSN 1058-5591.

Albert:XR-7-1-27

[Alb93b] Marc Albert. Best of netnews. The X Resource, 7(1):27–30, August 1993. CODEN XRESEA. ISBN 1-56592-022-8. ISSN 1058-5591.

Albert:XR-8-1-39

[Alb93c] Marc Albert. Best of netnews.

The X Resource, 8(1):39–43, October 1993. CODEN XRESEA. ISBN 1-56592-023-6. ISSN 1058-5591.

Albert:XR-10-1-55

[Alb94a] Marc Albert. Best of netnews.

The X Resource, 10(1):55–60, April 1994. CODEN XRESEA. ISBN 1-56592-067-8. ISSN 1058-5591.

Albert:XR-11-1-51

[Alb94b] Marc Albert. Best of netnews.

The X Resource, 11(1):51–56, July 1994. CODEN XRESEA. ISBN 1-56592-068-6. ISSN 1058-5591.

Albert:XR-12-1-51

[Alb94c] Marc Albert. Bet of netnews. The X Resource, 12(1):51–55, October 1994. CODEN XRESEA. ISBN 1-56592-069-4. ISSN 1058-5591.

Albert:XR-14-1-33

[Alb95a] Marc Albert. Best of netnews.

The X Resource, 14(1):33–36, April 1995. CODEN XRESEA. ISBN 1-56592-122-4. ISSN 1058-5591.

Albert:XR-15-1-35

[Alb95b] Marc Albert. Best of netnews.

The X Resource, 15(1):35–40, July 1995. CODEN XRESEA. ISBN 1-56592-140-2. ISSN 1058-5591.

Ackerman:XR-13-1-169

[AM95] Mark S. Ackerman and Eric Mandel. xbb: A blackboard for X applications. The X Resource, 13(1): 169–176, January 1995. CODEN XRESEA. ISBN 1-56592-121-6. ISSN 1058-5591.

Anonymous:XR-0-1-249

[Ano91a] Anonymous. How to publish in The X Resource. The X Resource, 0(1):249–253, October 1991. CO-DEN XRESEA. ISBN 0-937175-79-X. ISSN 1058-5591.

Anonymous:XR-0-1-37

[Ano91b] Anonymous. User group news. The X Resource, 0(1):37–38, October 1991. CODEN XRESEA. ISBN 0-937175-79-X. ISSN 1058-5591.

Anonymous:XR-4-1-164

[Ano92a] Anonymous. The CMAP widget. The X Resource, 4(1):164– 166, December 1992. CODEN XRESEA. ISBN 0-937175-99-4. ISSN 1058-5591.

Anonymous:XR-4-1-161

[Ano92b] Anonymous. A directory of freely and commercially available widgets. The X Resource, 4(1):161– 163, December 1992. CODEN XRESEA. ISBN 0-937175-99-4. ISSN 1058-5591.

Anonymous:XR-3-1-5

[Ano92c] Anonymous. From the editor. The X Resource, 3(1):5–6, June 1992. CODEN XRESEA. ISBN 0-937175-98-6. ISSN 1058-5591.

Anonymous:XR-4-1-167

[Ano92d] Anonymous. The Hdial widget.

The X Resource, 4(1):167–169, December 1992. CODEN XRE-SEA. ISBN 0-937175-99-4. ISSN 1058-5591.

Anonymous:XR-1-1-235

[Ano92e] Anonymous. How to publish in The X Resource. The X Resource, 1(1):235–239, January 1992. CO-DEN XRESEA. ISBN 0-937175-96-X. ISSN 1058-5591.

Anonymous:XR-3-1-7

[Ano92f] Anonymous. Obtaining the source code. The X Resource, 3(1):7–12, June 1992. CODEN XRESEA. ISBN 0-937175-98-6. ISSN 1058-5591.

Anonymous:XR-4-1-35

[Ano92g] Anonymous. X user groups. The X Resource, 4(1):35–38, December 1992. CODEN XRESEA. ISBN 0-937175-99-4. ISSN 1058-5591.

Anonymous:XR-7-1-9

[Ano93a] Anonymous. 8th annual X technical conference. The X Resource, 7 (1):9–10, August 1993. CODEN XRESEA. ISBN 1-56592-022-8. ISSN 1058-5591.

Anonymous:XR-8-1-9

[Ano93b] Anonymous. 8th annual X technical conference. The X Resource, 8 (1):9–10, October 1993. CODEN XRESEA. ISBN 1-56592-023-6. ISSN 1058-5591.

Anonymous:XR-6-1-227

[Ano93c] Anonymous. Back issues of the X Resource. The X Resource, 6(1): 227–232, March 1993. CODEN XRESEA. ISBN 1-56592-021-X. ISSN 1058-5591.

Anonymous:XR-6-1-5

[Ano93d] Anonymous. From the editor. The X Resource, 6(1):5–6, March 1993. CODEN XRESEA. ISBN 1-56592-021-X. ISSN 1058-5591.

Anonymous:XR-7-1-5

[Ano93e] Anonymous. From the editor. The X Resource, 7(1):5–6, August 1993. CODEN XRESEA. ISBN 1-56592-022-8. ISSN 1058-5591.

Anonymous:XR-8-1-5

[Ano93f] Anonymous. From the editor. The X Resource, 8(1):5–7, October 1993. CODEN XRESEA. ISBN 1-56592-023-6. ISSN 1058-5591.

Anonymous:XR-6-1-7

[Ano93g] Anonymous. Letters: RPC programming in X applications. The X Resource, 6(1):7–13, March 1993. CODEN XRESEA. ISBN 1-56592-021-X. ISSN 1058-5591.

Anonymous:XR-6-1-14

[Ano93h] Anonymous. Obtaining the source code. The X Resource, 6 (1):14–17, March 1993. CODEN XRESEA. ISBN 1-56592-021-X. ISSN 1058-5591.

Anonymous:XR-8-1-11

[Ano93i] Anonymous. Obtaining the source code. The X Resource, 8 (1):11–14, October 1993. CODEN XRESEA. ISBN 1-56592-023-6. ISSN 1058-5591.

Anonymous:XR-7-1-19

[Ano93j] Anonymous. The X consortium:

The next generation. The X Resource, 7(1):19–22, August 1993. CODEN XRESEA. ISBN 1-56592-022-8. ISSN 1058-5591.

Anonymous:XR-7-1-7

[Ano93k] Anonymous. The X Resource code archive. The X Resource, 7(1):7–8, August 1993. CODEN XRESEA. ISBN 1-56592-022-8. ISSN 1058-5591.

Anonymous:XR-12-1-7

[Ano94a] Anonymous. 9th annual technical conference: “The designers’ X conference”: Call for papers, talks and tutorials. The X Resource, 12 (1):7–8, October 1994. CODEN XRESEA. ISBN 1-56592-069-4. ISSN 1058-5591.

Anonymous:XR-9-1-5

[Ano94b] Anonymous. From the editor.

The X Resource, 9(1):5, January 1994. CODEN XRESEA. ISBN 1-56592-066-X. ISSN 1058-5591.

Anonymous:XR-10-1-7

[Ano94c] Anonymous. Obtaining the source code. The X Resource, 10(1):7–10, April 1994. CODEN XRESEA. ISBN 1-56592-067-8. ISSN 1058-5591.

Anonymous:XR-11-1-7

[Ano94d] Anonymous. Obtaining the source code. The X Resource, 11(1):7–10, July 1994. CODEN XRESEA. ISBN 1-56592-068-6. ISSN 1058-5591.

Anonymous:XR-12-1-9

[Ano94e] Anonymous. Obtaining the source code. The X Resource, 12 (1):9–12, October 1994. CODEN XRESEA. ISBN 1-56592-069-4. ISSN 1058-5591.

Asmus:XR-16-1-59

[Asm95] Madeline T. Asmus. The PanoramiX extension. The X Resource, 16(1):59–73, December 1995. CODEN XRESEA. ISBN 0-937175-79-X. ISSN 1058-5591.

Baeker:XR-1-2-106

[Bäc92]  Andreas Bäcker.  Designing

reusable widget classes with C++ and OSF/Motif. The X Resource, 2(1):106–130, April 1992. CO-DEN XRESEA. ISBN 0-937175-97-8. ISSN 1058-5591.

Ball:XR-5-1-167

[Bal93] Steven Ball. Distributed memory multicomputers as X clients. The X Resource, 5(1):167–181, January 1993. CODEN XRESEA. ISBN 1-56592-020-1. ISSN 1058-5591.

Bartlett:XR-1-1-117

[Bar92] Joel F. Bartlett. Don’t fidget with widgets, draw! The X Resource, 1(1):117–131, January 1992. CO-DEN XRESEA. ISBN 0-937175-96-X. ISSN 1058-5591.

Beged-Dov:XR-3-1-13

[BD92] Gabe Beged-Dov. Xt performance improvements in release 5. The X Resource, 3(1):13–24, June 1992. CODEN XRESEA. ISBN 0-937175-98-6. ISSN 1058-5591.

Beged-Dov:XR-1-1-169

[BDC92] Gabe Beged-Dov and Ellis S. Cohen. Implementing drag & drop for X11. The X Resource, 1(1): 169–190, January 1992. CODEN XRESEA. ISBN 0-937175-96-X. ISSN 1058-5591.

Benson:XR-10-1-167

[Ben94] Channing Benson. Implementing cut and paste in the X environment. The X Resource, 10(1):167– 198, April 1994. CODEN XRE-SEA. ISBN 1-56592-067-8. ISSN 1058-5591.

Berlage:XR-3-1-137

[Ber92] Thomas Berlage. A ruler model for the Motif form widget. The X Resource, 3(1):137–154, June 1992. CODEN XRESEA. ISBN 0-937175-98-6. ISSN 1058-5591.

Berry:XR-13-1-179

[Ber95] David T. Berry. Integrating A color management system with A Unix and X11 environment. The X Resource, 13(1):179–180, January 1995. CODEN XRESEA. ISBN 1-56592-121-6. ISSN 1058-5591.

Beust:XR-14-1-89

[Beu95] Cedric Beust. Koalatalk: An ICE-based light-weight message bus. The X Resource, 14(1):89– 107, April 1995. CODEN XRE-SEA. ISBN 1-56592-122-4. ISSN 1058-5591.

Bormann:XR-9-1-205

[BH94] Carsten Bormann and Gero Hoffman. Xmc and Xy — scalable window sharing and mobility, or from X protocol multiplexing to X protocol multicasting. The X Resource, 9(1):205–210, January 1994. CODEN XRESEA. ISBN 1-56592-066-X. ISSN 1058-5591.

Billman:XR-4-1-137

[BHM^{+}92] Richard Billman, Robert Hammond, Pat McElhatton, Ellen Nordahl Brandt, Francis Sung, and Nancy Yost. Workstation audio in the X environment. The X Resource, 4(1):137–158, December 1992. CODEN XRESEA. ISBN 0-937175-99-4. ISSN 1058-5591.

Bilow:XR-9-1-7

[Bil94] Steven Craig Bilow. Use cases, objects, and X: Improving ergonomics and usability through user centered design. The X Resource, 9(1):7–22, January 1994. CODEN XRESEA. ISBN 1-56592-066-X. ISSN 1058-5591.

Blewett:XR-8-1-101

[BK93] Doug Blewett and Meg Kilduff. xtent 3.0—improvements on programming with resources. The X Resource, 8(1):101–118, October 1993. CODEN XRESEA. ISBN 1-56592-023-6. ISSN 1058-5591.

Bloomer:XR-4-1-59

[Blo92] John Bloomer. RPC programming in X applications. The X Resource, 4(1):59–101, December 1992. CODEN XRESEA. ISBN 0-937175-99-4. ISSN 1058-5591.

Beck:XR-10-1-47

[BM94] Bob Beck and Art Mulder. The X administrator: Keeping start-up files simple in a mixed xdm/xinit environment. The X Resource, 10 (1):47–54, April 1994. CODEN XRESEA. ISBN 1-56592-067-8. ISSN 1058-5591.

Brenkosh:XR-11-1-91

[Bre94] Joseph P. Brenkosh. Developing an X protocol extension. The X Resource, 11(1):91–114, July 1994. CODEN XRESEA. ISBN 1-56592-068-6. ISSN 1058-5591.

Bashir:XR-7-1-87

[BS93] Imran Bashir and David M.

Sternlicht. A tale of two toolkits: Xt versus Interviews. The X Resource, 7(1):87–109, August 1993. CODEN XRESEA. ISBN 1-56592-022-8. ISSN 1058-5591.

Baird-Smith:XR-13-1-141

[BSK95] Anselm Baird-Smith and Philippe Kaplan. k-Edit: a communication library built on ICE. The X Resource, 13(1):141–149, January 1995. CODEN XRESEA. ISBN 1-56592-121-6. ISSN 1058-5591.

Buehler:XR-4-1-109

[Bue92] Kurt Buehler. The XGEN application generator: Constructing MOTIF GUIs for TTY programs. The X Resource, 4(1):109– 136, December 1992. CODEN XRESEA. ISBN 0-937175-99-4. ISSN 1058-5591.

Cameron:XR-0-1-81

[CCR91] Bruce M. Cameron, J. J. Camp, and R. A. Robb. Default colormap manipulation. The X Resource, 0(1):81–93, October 1991. CODEN XRESEA. ISBN 0-937175-79-X. ISSN 1058-5591.

Correia:XR-5-1-193

[CG93] Nuno M. Correia and Nuno M. Guimarães. TaX: A tool for building time dependent applications. The X Resource, 5(1):193–202, January 1993. CODEN XRE-SEA. ISBN 1-56592-020-1. ISSN 1058-5591.

Chakrabarti:XR-13-1-7

[Cha95] Sankar L. Chakrabarti. Testing X-Clients using Synlib and FocusMaps. The X Resource, 13 (1):7–21, January 1995. CODEN XRESEA. ISBN 1-56592-121-6. ISSN 1058-5591.

Chesler:XR-7-1-79

[Che93] Robert Chesler. Experiences in extending X to provide determinism. The X Resource, 7(1):79–86, August 1993. CODEN XRESEA. ISBN 1-56592-022-8. ISSN 1058-5591.

Cohen:XR-5-1-23

[Coh93] Ellis S. Cohen. Describing formats for X-based data interchange. The X Resource, 5(1):25– 52, January 1993. CODEN XRE-SEA. ISBN 1-56592-020-1. ISSN 1058-5591.

Connolly:XR-3-1-211

[Con92] Dan Connolly. Public-domain XcRichText widget. The X Resource, 3(1):211–216, June 1992. CODEN XRESEA. ISBN 0-937175-98-6. ISSN 1058-5591.

Courtney:XR-0-1-39

[Cou91] Antony A. Courtney. Engineering insights from an interactive imaging application. The X Resource, 0(1):39–59, October 1991. CO-DEN XRESEA. ISBN 0-937175-79-X. ISSN 1058-5591.

Daifuku:XR-5-1-239

[Dai93] Peter Daifuku. A fully functional implementation of layered windows. The X Resource, 5(1): 239–249, January 1993. CODEN XRESEA. ISBN 1-56592-020-1. ISSN 1058-5591.

Danskin:XR-13-1-23

[Dan95] John M. Danskin. Previewing PostScript over a telephone in 3 seconds per page. The X Resource, 13(1):23–40, January 1995. CODEN XRESEA. ISBN 1-56592-121-6. ISSN 1058-5591.

Dardailler:XR-1-1-143

[Dar92] Daniel Dardailler. Testing widget geometry management. The X Resource, 1(1):143–156, January 1992. CODEN XRESEA. ISBN 0-937175-96-X. ISSN 1058-5591.

Dardailler:XR-9-1-35

[Dar94] Daniel Dardailler. Traitifying Motif. The X Resource, 9(1):35– 46, January 1994. CODEN XRE-SEA. ISBN 1-56592-066-X. ISSN 1058-5591.

Dardailler:XR-15-1-71

[Dar95] Daniel Dardailler. X over the Web. The X Resource, 15(1):71– 83, July 1995. CODEN XRESEA. ISBN 1-56592-140-2. ISSN 1058-

Davey:XR-0-1-24

[Dav91] Paul Davey. The X user: xterm tips and tricks. The X Resource, 0 (1):24–30, October 1991. CODEN XRESEA. ISBN 0-937175-79-X. ISSN 1058-5591.

Davey:XR-1-2-35

[Dav92a] Paul Davey. The X user: Demystifying Imake. The X Resource, 2(1):35–48, April 1992. CODEN XRESEA. ISBN 0-937175-97-8. ISSN 1058-5591.

Davey:XR-3-1-25

[Dav92b] Paul Davey. The X user: Making better use of resources. The X Resource, 3(1):25–34, June 1992. CODEN XRESEA. ISBN 0-937175-98-6. ISSN 1058-5591.

Davey:XR-6-1-141

[Dav93a] Paul Davey. Three help systems: xhelp, X.deskhelp, and UNIXhelp. The X Resource, 6(1):141– 151, March 1993. CODEN XRE-SEA. ISBN 1-56592-021-X. ISSN 1058-5591.

Davey:XR-6-1-37

[Dav93b] Paul Davey. User-contributed software: Contrib: The other tree. The X Resource, 6(1):37–44, March 1993. CODEN XRESEA. ISBN 1-56592-021-X. ISSN 1058-5591.

Deininger:XR-10-1-141

[Dei94] Axel Deininger. An X print server: Bringing WYSIWYG to X. The X Resource, 10(1):141– 155, April 1994. CODEN XRE-SEA. ISBN 1-56592-067-8. ISSN 1058-5591.

Deininger:XR-3-1-97

[DM92] Axel Deininger and Nathan Meyers. Using the new font capabilities of HP-donated font server enhancements. The X Resource, 3 (1):97–120, June 1992. CODEN XRESEA. ISBN 0-937175-98-6. ISSN 1058-5591.

Danilevsky:XR-12-1-121

[DR94] Harry Danilevsky and Andrew Renalds. An object-oriented approach to Motif 1.2 drag and drop. The X Resource, 12(1): 121–131, October 1994. CODEN XRESEA. ISBN 1-56592-069-4. ISSN 1058-5591.

Dubois:XR-8-1-29

[Dub93] Paul Dubois. Imakefile troubleshooting. The X Resource, 8 (1):29–37, October 1993. CODEN XRESEA. ISBN 1-56592-023-6. ISSN 1058-5591.

Dykes:XR-0-1-215

[Dyk91] Gene Dykes. Xcu—Cornell University widget set. The X Resource, 0(1):215–248, October 1991. CODEN XRESEA. ISBN 0-937175-79-X. ISSN 1058-5591.

Dyksen:XR-1-2-78

[Dyk92] Wayne Dyksen. Controlling event delivery with grabs and keyboard focus. The X Resource, 2(1):78– 92, April 1992. CODEN XRE-SEA. ISBN 0-937175-97-8. ISSN 1058-5591.

Dykstra:XR-9-1-195

[Dyk94] Phillip Dykstra. X11 in virtual environments: Combining computer interaction methodologies.

The X Resource, 9(1):195–204, January 1994. CODEN XRE-SEA. ISBN 1-56592-066-X. ISSN 1058-5591.

Eliens:XR-14-1-111

[Eli95] Anton Eliëns. Hush: A C++ API for Tcl/Tk. The X Resource, 14 (1):111–155, April 1995. CODEN XRESEA. ISBN 1-56592-122-4. ISSN 1058-5591.

Elliott:XR-13-1-257

[Ell95a] Ian Elliott. The double buffer extension. The X Resource, 13 (1):257, January 1995. CODEN XRESEA. ISBN 1-56592-121-6. ISSN 1058-5591.

Elliott:XR-13-1-75

[Ell95b] Ian Elliott. Help! There’s a spy in my code. The X Resource, 13 (1):75–91, January 1995. CODEN XRESEA. ISBN 1-56592-121-6. ISSN 1058-5591.

Edwards:XR-13-1-245

[ELMW95] W. Keith Edwards, Susan H. Liebeskind, Elizabeth D. Mynatt, and William D. Walker. A remote access protocol for the X Window System. The X Resource, 13(1): 245–256, January 1995. CODEN XRESEA. ISBN 1-56592-121-6. ISSN 1058-5591.

Epstein:XR-0-1-113

[Eps91] Jeremy Epstein. Issues in building trusted X Window Systems. The X Resource, 0(1):113–132, October 1991. CODEN XRESEA. ISBN 0-937175-79-X. ISSN 1058-5591.

Epstein:XR-4-1-103

[Eps92] Jeremy Epstein. A bibliography of windowing systems and security. The X Resource, 4(1):103– 108, December 1992. CODEN XRESEA. ISBN 0-937175-99-4. ISSN 1058-5591.

Edwards:XR-5-1-229

[ER93] W. Keith Edwards and Tom Rodriguez. Runtime translation of X interfaces to support visuallyimpaired users. The X Resource, 5(1):229–238, January 1993. CO-DEN XRESEA. ISBN 1-56592-020-1. ISSN 1058-5591.

Evans:XR-13-1-217

[Eva95] Steve Evans. Common Desktop Environment architectural overview. The X Resource, 13(1): 217–234, January 1995. CODEN XRESEA. ISBN 1-56592-121-6. ISSN 1058-5591.

Fahy:XR-9-1-105

[Fah94] J. Ben Fahy. Experience with XIE: Server and client, past, present and future. The X Resource, 9(1):105–120, January 1994. CODEN XRESEA. ISBN 1-56592-066-X. ISSN 1058-5591.

Ferguson:XR-4-1-171

[Fer92a] Paula Ferguson. The X11 input extension: A tutorial. The X Resource, 4(1):171–194, December 1992. CODEN XRESEA. ISBN 0-937175-99-4. ISSN 1058-5591.

Ferguson:XR-4-1-195

[Fer92b] Paula Ferguson. The X11 input extension: Reference pages. The X Resource, 4(1):195–270, December 1992. CODEN XRESEA. ISBN 0-937175-99-4. ISSN 1058-5591.

Ferguson:XR-3-1-179

[Fer92c] Paula M. Ferguson. The nonrectangular window shape extension. The X Resource, 3(1):179– 210, June 1992. CODEN XRE-SEA. ISBN 0-937175-98-6. ISSN 1058-5591.

Ferguson:XR-11-1-6

[Fer94a] Paula M. Ferguson. From the editor. The X Resource, 11(1):5– 6, July 1994. CODEN XRESEA. ISBN 1-56592-068-6. ISSN 1058-5591.

Ferguson:XR-12-1-5

[Fer94b] Paula M. Ferguson. From the editor. The X Resource, 12(1):5– 6, October 1994. CODEN XRE-SEA. ISBN 1-56592-069-4. ISSN 1058-5591.

Ferguson:XR-16-1-7

[Fer95a] Paula M. Ferguson. 10th Annual X Technical Conference. The X Resource, 16(1):7–11, December 1995. CODEN XRESEA. ISBN 0-937175-79-X. ISSN 1058-5591.

Ferguson:XR-13-1-5

[Fer95b] Paula M. Ferguson. From the editor. The X Resource, 13(1):5, January 1995. CODEN XRESEA. ISBN 1-56592-121-6. ISSN 1058-5591.

Ferguson:XR-14-1-5

[Fer95c] Paula M. Ferguson. From the editor. The X Resource, 14(1):5–6, April 1995. CODEN XRESEA. ISBN 1-56592-122-4. ISSN 1058-5591.

Ferguson:XR-16-1-5

[Fer95d] Paula M. Ferguson. From the editor. The X Resource, 16(1):5–6, December 1995. CODEN XRE-SEA. ISBN 0-937175-79-X. ISSN 1058-5591.

Ferguson:XR-14-1-7

[Fer95e] Paula M. Ferguson. Obtaining the source code. The X Resource, 14(1):7–10, April 1995. CODEN XRESEA. ISBN 1-56592-122-4. ISSN 1058-5591.

Ferguson:XR-16-1-13

[Fer95f] Paula M. Ferguson. Obtaining the source code. The X Resource, 16 (1):13–16, December 1995. CO-DEN XRESEA. ISBN 0-937175-79-X. ISSN 1058-5591.

Fulton:XR-5-1-251

[FK93] Jim Fulton and Chris Kent Kantarjiev. An update on low bandwidth X (LBX). The X Resource, 5(1):251–266, January 1993. CO-DEN XRESEA. ISBN 1-56592-020-1. ISSN 1058-5591.

Flanagan:XR-4-1-45

[Fla92] David Flanagan. Finding auxiliary files for an X toolkit application. The X Resource, 4(1):45–57, December 1992. CODEN XRE-SEA. ISBN 0-937175-99-4. ISSN 1058-5591.

Flanagan:XR-8-1-87

[Fla93a] David Flanagan. Widget architecture for data display. The X Resource, 8(1):87–100, October 1993. CODEN XRESEA. ISBN 1-56592-023-6. ISSN 1058-5591.

Flanagan:XR-8-1-17

[Fla93b] David Flanagan. The X user— drawing figures with xfig. The X Resource, 8(1):17–27, October 1993. CODEN XRESEA. ISBN 1-56592-023-6. ISSN 1058-5591.

Flanagan:XR-11-1-173

[Fla94] David Flanagan. The XmtLayout widget: A tutorial. The X Resource, 11(1):173–204, July 1994. CODEN XRESEA. ISBN 1-56592-068-6. ISSN 1058-5591.

Flowers:XR-D-1-G

[Flo94] Jim Flowers. X logical font description conventions: Version 1.5: Public review draft (with change bars against version 1.4): X consortium draft standard. The X Resource, Special issue D(1): G–i–G–v, G–1–G–28, February 1994. CODEN XRESEA. ISBN 1-56592-071-6. ISSN 1058-5591.

Fortune:XR-9-1-135

[For94] Erik Fortune. The X keyboard (XKB) extension: Overview and rationale. The X Resource, 9(1): 135–149, January 1994. CODEN XRESEA. ISBN 1-56592-066-X. ISSN 1058-5591.

Fulton:XR-9-1-181

[FR94] Jim Fulton and Greg Renda. The network audio system: Make your applications sing (as well as dance)! The X Resource, 9(1): 181–194, January 1994. CODEN XRESEA. ISBN 1-56592-066-X. ISSN 1058-5591.

Freedman:XR-14-1-13

[Fre95] Matt Freedman. The X user:

WILLOW—the Washington Information Looker-Upper Layered Over Windows. The X Resource, 14(1):13–31, April 1995. CODEN XRESEA. ISBN 1-56592-122-4. ISSN 1058-5591.

Gaskins:XR-B-1-1a

[Gas92] Tom Gaskins. Introduction to PEXlib. The X Resource, Special issue B(1):1–20, September 1992. CODEN XRESEA. ISBN 1-56592-050-3. ISSN 1058-5591.

Gaskins:XR-8-1-121

[Gas93] Tom Gaskins. The multi-buffering extension—smooth animation effortlessly. The X Resource, 8(1): 121–159, October 1993. CODEN XRESEA. ISBN 1-56592-023-6. ISSN 1058-5591.

Glauert:XR-D-1-I

[GCGW94] Tim Glauert, Dave Carver, Jim Gettys, and David P. Wiggins. X synchronization extension: Public review draft: X consortium draft standard. The X Resource, Special issue D(1):H–i–H–iv, H– 1–H–37, February 1994. CODEN XRESEA. ISBN 1-56592-071-6. ISSN 1058-5591.

Gemeinhart:XR-16-1-19

[Gem95] Keith Gemeinhart. An elegant online help system for X/Motif programs. The X Resource, 16(1): 19–30, December 1995. CODEN XRESEA. ISBN 0-937175-79-X. ISSN 1058-5591.

Gendler:XR-7-1-113

[Gen93] Bruce Gendler. The Widget Method statistics package: Tools for monitoring and measuring method usage. The X Resource, 7(1):113–133, August 1993. CO-DEN XRESEA. ISBN 1-56592-022-8. ISSN 1058-5591.

Gildea:XR-4-1-15

[Gil92] Stephen Gildea. The MIT X software distribution after release 5. The X Resource, 4(1):15–18, December 1992. CODEN XRESEA. ISBN 0-937175-99-4. ISSN 1058-5591.

Gildea:XR-5-1-159

[Gil93] Stephen Gildea. Multi-threaded Xlib. The X Resource, 5(1): 159–166, January 1993. CODEN XRESEA. ISBN 1-56592-020-1. ISSN 1058-5591.

Gildea:XR-11-1-13

[Gil94] Stephen Gildea. From the X consortium: Writing portable X code. The X Resource, 11(1):13– 25, July 1994. CODEN XRESEA. ISBN 1-56592-068-6. ISSN 1058-5591.

Gourlay:XR-6-1-51

[Gou93] Alastair Gourlay. Writing Motif widgets. The X Resource, 6 (1):51–78, March 1993. CODEN XRESEA. ISBN 1-56592-021-X. ISSN 1058-5591.

Gourlay:XR-10-1-63

[Gou94] Alastair Gourlay. The one-minute manager: Custom Motif layout widgets made easy. The X Resource, 10(1):63–90, April 1994.

CODEN XRESEA. ISBN 1-56592-067-8. ISSN 1058-5591.

Grunwald:XR-0-1-184

[GT91] Dirk Grunwald and Bryan Totty.

Free Widget Foundation. The X Resource, 0(1):184–205, October 1991. CODEN XRESEA. ISBN 0-937175-79-X. ISSN 1058-5591.

Harrison:XR-0-1-206

[Har91] David Harrison. Wcl table widget. The X Resource, 0(1):206– 214, October 1991. CODEN XRESEA. ISBN 0-937175-79-X. ISSN 1058-5591.

Harris:XR-4-1-39

[Har92] Kathy Harris. The PEX interoperability center. The X Resource, 4(1):39–41, December 1992. CO-DEN XRESEA. ISBN 0-937175-99-4. ISSN 1058-5591.

Hemmer:XR-6-1-153

[Hem93] James J. Hemmer. Xhyper: A hypertext online help system. The X Resource, 6(1):153–164, March 1993. CODEN XRESEA. ISBN 1-56592-021-X. ISSN 1058-5591.

Hesketh:XR-12-1-21

[Hes94] Richard Hesketh. The X administrator: Configuration and administration of a scalable X-based UNIX service. The X Resource, 12(1):21–49, October 1994. CO-DEN XRESEA. ISBN 1-56592-069-4. ISSN 1058-5591.

Hare:XR-13-1-183

[HS95] Larry Hare and Bob Shelley. PhotoKit: An XIE toolkit. The X Resource, 13(1):183–199, January 1995. CODEN XRESEA. ISBN 1-56592-121-6. ISSN 1058-5591.

Humphrey:XR-5-1-53

[Hum93] Steve Humphrey. A widget class extension for improved geometry management. The X Resource, 5 (1):53–66, January 1993. CODEN XRESEA. ISBN 1-56592-020-1. ISSN 1058-5591.

Jacobi:XR-1-1-157

[Jac92] Christian P. Jacobi. Migrating widgets. The X Resource, 1(1): 157–167, January 1992. CODEN XRESEA. ISBN 0-937175-96-X. ISSN 1058-5591.

Jones:XR-3-1-55

[Jon92] Oliver Jones. Multi-user application software using Xt. The X Resource, 3(1):55–76, June 1992. CODEN XRESEA. ISBN 0-937175-98-6. ISSN 1058-5591.

Jones:XR-6-1-97

[Jon93] Oliver Jones. Multidisplay software in X: A survey of architectures. The X Resource, 6(1):97– 113, March 1993. CODEN XRE-SEA. ISBN 1-56592-021-X. ISSN 1058-5591.

Jones:XR-9-1-91

[Jon94] Peter C. Jones. Xvan: A true multiple screen X server. The X Resource, 9(1):91–104, January 1994. CODEN XRESEA. ISBN 1-56592-066-X. ISSN 1058-5591.

Karlton:XR-1-1-27

[Kar92] Phil Karlton. Integrating the GL into the X environment: A high performance rendering extension working with and not against X. The X Resource, 1(1):27–32, January 1992. CODEN XRESEA. ISBN 0-937175-96-X. ISSN 1058-5591.

Kempf:XR-13-1-151

[KBSW95] James Kempf, Sue Booker, Ethan Solomita, and Jackson Wong. Re-engineering the CDE/Motif graphical user interface for pens. The X Resource, 13(1):151–168, January 1995. CODEN XRE-SEA. ISBN 1-56592-121-6. ISSN 1058-5591.

Kempf:XR-11-1-117

[Kem94] James Kempf. Pseudoservers: A survey. The X Resource, 11(1): 117–125, July 1994. CODEN XRESEA. ISBN 1-56592-068-6. ISSN 1058-5591.

Kepple:XR-3-1-155

[Kep92] Lawrence R. Kepple. A new paradigm for cross-platform automated GUI testing. The X Resource, 3(1):155–178, June 1992. CODEN XRESEA. ISBN 0-937175-98-6. ISSN 1058-5591.

Kilgard:XR-9-1-73

[KHLS94] Mark J. Kilgard, Simon Hui, Allen A. Leinwand, and Dave Spalding. X server multirendering for OpenGL and PEX. The X Resource, 9(1):73–88, January 1994. CODEN XRESEA. ISBN 1-56592-066-X. ISSN 1058-5591.

Kilgard:XR-3-1-77

[Kil92] Mark J. Kilgard. An extension for server instrumentation and tracing. The X Resource, 3(1):77–96, June 1992. CODEN XRESEA. ISBN 0-937175-98-6. ISSN 1058-5591.

Kilgard:XR-13-1-41

[Kil95] Mark J. Kilgard. D11: a highperformance protocol-optional, transport-optional window system with X11 compatibility and semantics. The X Resource, 13 (1):41–61, January 1995. CODEN XRESEA. ISBN 1-56592-121-6. ISSN 1058-5591.

Kempf:XR-5-1-203

[KW93] James Kempf and Alan Wilson. Supporting mobile, penbased computing with X. The X Resource, 5(1):203–211, January 1993. CODEN XRESEA. ISBN 1-56592-020-1. ISSN 1058-5591.

Lazear:XR-16-1-45

[Laz95] Walter D. Lazear. Visualizing X Window System performance. The X Resource, 16(1):45–57, December 1995. CODEN XRESEA. ISBN 0-937175-79-X. ISSN 1058-5591.

LeHors:XR-12-1-135

[Le 94] Arnaud Le Hors. The XPM format and library: A tutorial. The X Resource, 12(1):135–206, October 1994. CODEN XRESEA. ISBN 1-56592-069-4. ISSN 1058-5591.

Lembree:XR-13-1-259

[LE95] Rob Lembree and Ian Elliott. Loadable server architectures. The X Resource, 13(1):259–281, January 1995. CODEN XRE-SEA. ISBN 1-56592-121-6. ISSN 1058-5591.

Leon:XR-14-1-39

[Léo95] Jean-Michel Léon. The Knvas widget: A 2D graphics framework. The X Resource, 14(1):39– 58, April 1995. CODEN XRE-SEA. ISBN 1-56592-122-4. ISSN 1058-5591.

Lin:XR-1-1-191

[Lin92a] Jin-Kun Lin. Virtual screen: A framework for task management. The X Resource, 1(1):191–198, January 1992. CODEN XRE-SEA. ISBN 0-937175-96-X. ISSN 1058-5591.

Linton:XR-1-1-109

[Lin92b] Mark A. Linton. Implementing resolution independence on top of the X Window System. The X Resource, 1(1):109–116, January 1992. CODEN XRESEA. ISBN 0-937175-96-X. ISSN 1058-5591.

Linton:XR-5-1-77

[LP93] Mark Linton and Chuck Price.

Building distributed user interfaces with Fresco. The X Resource, 5(1):77–87, January 1993. CODEN XRESEA. ISBN 1-56592-020-1. ISSN 1058-5591.

Lemke:XR-1-2-131

[LR92] David Lemke and David Rosenthal. Visualizing X. The X Resource, 2(1):131–153, April 1992. CODEN XRESEA. ISBN 0-937175-97-8. ISSN 1058-5591.

Linton:XR-9-1-63

[LTC94] Mark Linton, Steven Tang, and Steven Churchill. Redisplay in Fresco. The X Resource, 9(1):63– 69, January 1994. CODEN XRE-SEA. ISBN 1-56592-066-X. ISSN 1058-5591.

Main:XR-9-1-167

[Mai94] F. Beachley Main. The POSC/ Halliburton shared colormap system: Distributed management of X color resources. The X Resource, 9(1):167–179, January 1994. CODEN XRESEA. ISBN 1-56592-066-X. ISSN 1058-5591.

Malacarne:XR-1-1-7

[Mal92] Jeff Malacarne. Can you bet your life on X? Using the X Window System for command and control displays. The X Resource, 1(1):7– 25, January 1992. CODEN XRE-SEA. ISBN 0-937175-96-X. ISSN 1058-5591.

Manasse:XR-5-1-107

[Man93] Mark S. Manasse. The Trestle Toolkit. The X Resource, 5(1): 107–112, January 1993. CODEN XRESEA. ISBN 1-56592-020-1. ISSN 1058-5591.

Mark:XR-11-1-205

[Mar94a] Graham A. Mark. A tutorial introduction to Tcl and Tk. The X Resource, 11(1):205–248, July 1994. CODEN XRESEA. ISBN 1-56592-068-6. ISSN 1058-5591.

Marks:XR-9-1-155

[Mar94b] Stuart W. Marks. Inter-client communication in X11R6 and beyond. The X Resource, 9(1): 155–161, January 1994. CODEN XRESEA. ISBN 1-56592-066-X. ISSN 1058-5591.

McCormack:XR-D-1-A

[MAS94] Joel McCormack, Paul Asente, and Ralph R. Swick. X toolkit intrinsics — C language interface: X window system: X version 11, release 6: Public review draft: X consortium draft standard. The X Resource, Special issue D(1):A–i–A–x, A–1–A–71, February 1994. CODEN XRE-SEA. ISBN 1-56592-071-6. ISSN 1058-5591.

Marks:XR-5-1-149

[MC93] Stuart W. Marks and Laurence P. G. Cable. Trace analysis of the X Window System protocol. The X Resource, 5(1):149–158, January 1993. CODEN XRESEA. ISBN 1-56592-020-1. ISSN 1058-5591.

McGregor:XR-0-1-95

[McG91] Scott McGregor. Prescient agents: A Radar O’Reilly for your desktop. The X Resource, 0(1): 95–111, October 1991. CODEN XRESEA. ISBN 0-937175-79-X. ISSN 1058-5591.

McGregor:XR-1-2-93

[McG92] Scott McGregor. Designing for usability. The X Resource, 2(1):93– 105, April 1992. CODEN XRE-SEA. ISBN 0-937175-97-8. ISSN 1058-5591.

Menges:XR-5-1-129

[Men93] John Menges. The X engine library: A C++ library for constructing X pseudo-servers. The X Resource, 5(1):129–141, January 1993. CODEN XRESEA. ISBN 1-56592-020-1. ISSN 1058-5591.

Meyers:XR-9-1-211

[Mey94] Nathan Meyers. New font technology for X11R6: Changes for R6 increase font rendering capabilities, improve performance, and enable licensing. The X Resource, 9(1):211–227, January 1994. CODEN XRESEA. ISBN 1-56592-066-X. ISSN 1058-5591.

McCormack:XR-9-1-89

[MMG94] Joel McCormack, Bob McNamara, and Chris Gianos. The smart frame buffer goes Hollywood. The X Resource, 9(1):89– 90, January 1994. CODEN XRE-SEA. ISBN 1-56592-066-X. ISSN 1058-5591.

McNutt:XR-1-2-14

[MO92] Dinah McNutt and Miles O’Neal. The X administrator: Font formats and utilities. The X Resource, 2(1):14–34, April 1992. CODEN XRESEA. ISBN 0-937175-97-8. ISSN 1058-5591.

Mor:XR-6-1-21

[Mor93] Ralph Mor. From the X Consortium: PEXlib: A new 3-D graphics API for the X window system. The X Resource, 6(1):21–25, March 1993. CODEN XRESEA. ISBN 1-56592-021-X. ISSN 1058-5591.

Mor:XR-D-1-C

[Mor94a] Ralph Mor. Inter-client exchange (ICE) library: Public review draft: X consortium draft standard. The X Resource, Special issue D(1):C–i–C–iii, C–1–C–28, February 1994. CODEN XRE-SEA. ISBN 1-56592-071-6. ISSN 1058-5591.

Mor:XR-D-1-E

[Mor94b] Ralph Mor. X session management library: Public review draft: X consortium draft standard. The X Resource, Special issue D(1):E– i–E–iv, E–1–E–27, February 1994. CODEN XRESEA. ISBN 1-56592-071-6. ISSN 1058-5591.

Mandel:XR-13-1-235

[MST95] Eric Mandel, Ralph Swick, and Doug Tody. The X public access mechanism: Software cooperation for space science and beyond. The X Resource, 13(1):235–244, January 1995. CODEN XRESEA. ISBN 1-56592-121-6. ISSN 1058-5591.

Nath:XR-9-1-23

[Nat94] Satyajit Nath. Zero-defect widgets: A formal approach to testing Xt-based toolkits. The X Resource, 9(1):23–33, January 1994. CODEN XRESEA. ISBN 1-56592-066-X. ISSN 1058-5591.

Nelson:XR-13-1-107

[Nel95] Chris Nelson. OpenDoc and its architecture. The X Resource, 13 (1):107–126, January 1995. CO-DEN XRESEA. ISBN 1-56592-121-6. ISSN 1058-5591.

Newman:XR-1-1-49

[New92] Todd Newman. How not to implement overlays in X. The X Resource, 1(1):49–60, January 1992. CODEN XRESEA. ISBN 0-937175-96-X. ISSN 1058-5591.

Newmarch:XR-13-1-93

[New95a] Jan Newmarch. Embedding of X applications. The X Resource, 13 (1):93–105, January 1995. CO-DEN XRESEA. ISBN 1-56592-121-6. ISSN 1058-5591.

Newmarch:XR-15-1-19

[New95b] Jan Newmarch. The X user: XScript—shell programming with X. The X Resource, 15(1):19–34, July 1995. CODEN XRESEA. ISBN 1-56592-140-2. ISSN 1058-5591.

Narita:XR-D-1-H

[NH94] Masahiko Narita and Hideki Hiura. The input method protocol: Public review draft: X consortium draft standard. The X Resource, Special issue D(1): H–i–H–iv, H–1–H–37, February 1994. CODEN XRESEA. ISBN 1-56592-071-6. ISSN 1058-5591.

Niemeyer:XR-16-1-77

[Nie95] Pat Niemeyer. Exploring Java: A tutorial introduction. The X Resource, 16(1):77–102, December 1995. CODEN XRESEA. ISBN 0-937175-79-X. ISSN 1058-5591.

Nisewanger:XR-5-1-143

[Nis93] Jeff Nisewanger. The X file system. The X Resource, 5(1): 143–148, January 1993. CODEN XRESEA. ISBN 1-56592-020-1. ISSN 1058-5591.

Neville-Neil:XR-8-1-69

[NN93] George V. Neville-Neil. Current efforts in Client/Server audio. The X Resource, 8(1):69– 86, October 1993. CODEN XRE-SEA. ISBN 1-56592-023-6. ISSN 1058-5591.

Netousek:XR-11-1-127

[NP94] Thomas K. Netousek and Christoph G.

Prinz. A window watcher for X: Control of X clients by spoken words. The X Resource, 11 (1):127–142, July 1994. CODEN XRESEA. ISBN 1-56592-068-6. ISSN 1058-5591.

Masahiko:XR-10-1-125

[NS94] Masahiko Narita and Makiko Shimamura. Overview of the input method protocol. The X Resource, 10(1):125–139, April 1994. CODEN XRESEA. ISBN 1-56592-067-8. ISSN 1058-5591.

Nye:XR-0-1-31

[Nye91] Adrian Nye. The X administrator: Building X11R5 in limited disk space. The X Resource, 0 (1):31–35, October 1991. CODEN XRESEA. ISBN 0-937175-79-X. ISSN 1058-5591.

Nye:XR-7-1-65

[Nye93] Adrian Nye. X application debugging. The X Resource, 7(1):65–78, August 1993. CODEN XRESEA. ISBN 1-56592-022-8. ISSN 1058-5591.

Nye:XR-10-1-5

[Nye94] Adrian Nye. From the editor.

The X Resource, 10(1):5–6, April 1994. CODEN XRESEA. ISBN 1-56592-067-8. ISSN 1058-5591.

Nye:XR-15-1-5

[Nye95a] Adrian Nye. From the editor.

The X Resource, 15(1):5–6, July 1995. CODEN XRESEA. ISBN 1-56592-140-2. ISSN 1058-5591.

Nye:XR-15-1-7

[Nye95b] Adrian Nye. Motif ’95 and CDE developers conference. The X Resource, 15(1):7–11, July 1995. CODEN XRESEA. ISBN 1-56592-140-2. ISSN 1058-5591.

Nye:XR-15-1-13

[Nye95c] Adrian Nye. Obtaining the source code. The X Resource, 15(1):13– 16, July 1995. CODEN XRESEA. ISBN 1-56592-140-2. ISSN 1058-5591.

ONeal:XR-4-1-19

[O’N92] Miles O’Neal. The X administrator: Taming the X display manager. The X Resource, 4(1):19–34, December 1992. CODEN XRE-SEA. ISBN 0-937175-99-4. ISSN 1058-5591.

ONeal:XR-7-1-23

[O’N93] Miles O’Neal. The X user: xrdb and screen-specific resources. The X Resource, 7(1):23–26, August 1993. CODEN XRESEA. ISBN 1-56592-022-8. ISSN 1058-5591.

Ousterhout:XR-5-1-113

[Ous93] John Ousterhout. Hypergraphics and hypertext in Tk. The X Resource, 5(1):113–127, January 1993. CODEN XRESEA. ISBN 1-56592-020-1. ISSN 1058-5591.

Packard:XR-1-1-199

[Pac92] Keith Packard. Using XTrap to help people with manual disabilities. The X Resource, 1(1): 199–211, January 1992. CODEN XRESEA. ISBN 0-937175-96-X. ISSN 1058-5591.

Packard:XR-5-1-67

[Pac93] Keith Packard. The layout widget: A hierarchical constraint based widget. The X Resource, 5 (1):67–76, January 1993. CODEN XRESEA. ISBN 1-56592-020-1. ISSN 1058-5591.

Packard:XR-9-1-121

[Pac94] Keith Packard. Designing LBX:

An experiment based standard. The X Resource, 9(1):121–133, January 1994. CODEN XRE-SEA. ISBN 1-56592-066-X. ISSN 1058-5591.

Packard:XR-13-1-127

[Pac95] Keith Packard. A pseudo-root extension: X Window System nesting on a budget. The X Resource, 13(1):127–132, January 1995. CODEN XRESEA. ISBN 1-56592-121-6. ISSN 1058-5591.

Parisi:XR-0-1-71

[Par91] Tony Parisi. X and MS-Windows:

Notes on portability. The X Resource, 0(1):71–80, October 1991. CODEN XRESEA. ISBN 0-937175-79-X. ISSN 1058-5591.

Peterson:XR-3-1-121

[PC92] Chris D. Peterson and Sharon E.

Chang. Improving X application performance. The X Resource, 3 (1):121–136, June 1992. CODEN XRESEA. ISBN 0-937175-98-6. ISSN 1058-5591.

Pearce:XR-3-1-35

[Pea92] Eric Pearce. The X administrator: Managing font servers.

The X Resource, 3(1):35–46, June 1992. CODEN XRESEA. ISBN 0-937175-98-6. ISSN 1058-5591.

Peterson:XR-0-1-154

[Pet91a] Chris D. Peterson. editres— A graphical resource editor for users and programmers. The X Resource, 0(1):154–171, October 1991. CODEN XRESEA. ISBN 0-937175-79-X. ISSN 1058-5591.

Peterson:XR-0-1-172

[Pet91b] Chris D. Peterson. editres— manpage. The X Resource, 0(1): 172–178, October 1991. CODEN XRESEA. ISBN 0-937175-79-X. ISSN 1058-5591.

Peterson:XR-0-1-179

[Pet91c] Chris D. Peterson. editres— protocol. The X Resource, 0(1): 179–183, October 1991. CODEN XRESEA. ISBN 0-937175-79-X. ISSN 1058-5591.

Phelps:XR-10-1-33

[Phe94] Thomas A. Phelps. The X user: TkMan—A man born again. The X Resource, 10(1):33–46, April 1994. CODEN XRESEA. ISBN 1-56592-067-8. ISSN 1058-5591.

Pinkerton:XR-12-1-93

[Pin94]  Glenn Pinkerton.  Event-free

structured graphics in the X environment. The X Resource, 12(1): 93–119, October 1994. CODEN XRESEA. ISBN 1-56592-069-4. ISSN 1058-5591.

Price:XR-9-1-47

[Pri94] Chuck Price. Extending Xt to support CORBA-based embedding. The X Resource, 9(1):47– 61, January 1994. CODEN XRE-SEA. ISBN 1-56592-066-X. ISSN 1058-5591.

Punkka:XR-13-1-201

[Pun95]  Veikko Punkka.  Intelligent

reusable user interface components. The X Resource, 13(1): 201–215, January 1995. CODEN XRESEA. ISBN 1-56592-121-6. ISSN 1058-5591.

Pong:XR-10-1-113

[PZL94] Man-Chi Pong, Youngguang Zhang, and Pak-Keung Lai. The HBF font format: Optimizing fixed-pitch font support. The X Resource, 10(1):113–123, April 1994. CODEN XRESEA. ISBN 1-56592-067-8. ISSN 1058-5591.

Quercia:XR-11-1-27

[Que94] Valerie Quercia. The X user: The fvwm window manager. The X Resource, 11(1):27–50, July 1994. CODEN XRESEA. ISBN 1-56592-068-6. ISSN 1058-5591.

Raney:XR-11-1-59

[Ran94] Scott Raney. Interactive GUI development environments: A comparison of Tcl/Tk, the desktop Kornshell, and MetaCard. The X Resource, 11(1):59–89, July 1994. CODEN XRESEA. ISBN 1-56592-068-6. ISSN 1058-5591.

Rand:XR-16-1-31

[Ran95] Douglas S. Rand. GoldenGate:

Automatic conversion and the MediaWarehouse. The X Resource, 16(1):31–44, December 1995. CODEN XRESEA. ISBN 0-937175-79-X. ISSN 1058-5591.

Raves:XR-1-1-33

[Rav92] William Raves. A PostScript X server. The X Resource, 1(1):33– 47, January 1992. CODEN XRE-SEA. ISBN 0-937175-96-X. ISSN 1058-5591.

Rawal:XR-1-1-133

[Raw92] Kuntal Rawal. A macro facility for X: Object specific event simulation. The X Resource, 1(1): 133–142, January 1992. CODEN XRESEA. ISBN 0-937175-96-X. ISSN 1058-5591.

Rand:XR-5-1-89

[RB93] Douglas S. Rand and Gilles Benati. Melding OSF/Motif, C++ and the intrinsics. The X Resource, 5(1):89–106, January 1993. CODEN XRESEA. ISBN 1-56592-020-1. ISSN 1058-5591.

Rennison:XR-1-1-213

[RBKL92] Earl Rennison, Rusti Baker, Doohyun David Kim, and Young-Hwan Lim. MuX: An X coexistent, time-based multimedia I/O server. The X Resource, 1(1): 213–233, January 1992. CODEN XRESEA. ISBN 0-937175-96-X. ISSN 1058-5591.

S:XR-10-1-157

[RC94a] Douglas S. Rand and Ellis Cohen. Automatic pop-up menus in Motif. The X Resource, 10(1):157– 163, April 1994. CODEN XRE-SEA. ISBN 1-56592-067-8. ISSN 1058-5591.

S:XR-10-1-91

[RC94b] Douglas S. Rand and Ellis Cohen. The Motif 2.0 uniform transfer model: Unifying selection, clipboard and drag and drop. The X Resource, 10(1):91–111, April 1994. CODEN XRESEA. ISBN 1-56592-067-8. ISSN 1058-5591.

Rechsteiner:XR-6-1-183

[RG93] Hubert Rechsteiner and Fran cois Grize. The font handler package. The X Resource, 6(1):183– 216, March 1993. CODEN XRE-SEA. ISBN 1-56592-021-X. ISSN 1058-5591.

Richardson:XR-13-1-133

[Ric95] Tristan Richardson. Teleporting: Mobile X sessions. The X Resource, 13(1):133–140, January 1995. CODEN XRESEA. ISBN 1-56592-121-6. ISSN 1058-5591.

Rosenthal:XR-D-1-F

[RM94] David Rosenthal and Stuart W.

Marks. Inter-client communication conventions manual: Version 2.0: Public review draft (with change bars against version 1.1): X consortium draft standard. The X Resource, Special issue D(1): F–i–F–viii, F–1–F–61, February 1994. CODEN XRESEA. ISBN 1-56592-071-6. ISSN 1058-5591.

Richards:XR-7-1-33

[RMR93] W. Keith Richards, Elizabeth D.

Mynatt, and Tom Rodriguez. The Mercator Project: A nonvisual interface to the X window system. The X Resource, 7(1):33–53, August 1993. CODEN XRESEA. ISBN 1-56592-022-8. ISSN 1058-

Ross:XR-7-1-55

[Ros93] George D. M. Ross. Managing X in a large distributed environment. The X Resource, 7(1):55– 64, August 1993. CODEN XRE-SEA. ISBN 1-56592-022-8. ISSN 1058-5591.

Scheifler:XR-D-1-B

[SB94] Robert Scheifler and Jordan Brown. Inter-client exchange (ICE) proposal: Public review draft: X consortium draft standard. The X Resource, Special issue D(1):B–i–B–iii, B–1–B–15, February 1994. CODEN XRE-SEA. ISBN 1-56592-071-6. ISSN 1058-5591.

Scheifler:XR-0-1-7

[Sch91] Bob Scheifler. From the X Consortium: What’s new in Release 5. The X Resource, 0(1):7–15, October 1991. CODEN XRESEA. ISBN 0-937175-79-X. ISSN 1058-5591.

Scheifler:XR-1-2-8

[Sch92] Bob Scheifler. From the X Consortium: Current projects. The X Resource, 2(1):8–13, April 1992. CODEN XRESEA. ISBN 0-937175-97-8. ISSN 1058-5591.

Scheifler:XR-7-1-13

[Sch93a] Robert Scheifler. From the X consortium: Buddy, can you spare an RPC? The X Resource, 7(1):13– 17, August 1993. CODEN XRE-SEA. ISBN 1-56592-022-8. ISSN 1058-5591.

Schmidgall:XR-6-1-127

[Sch93b] Jay D. Schmidgall. Implementing context help in Motif. The X Resource, 6(1):127–139, March 1993. CODEN XRESEA. ISBN 1-56592-021-X. ISSN 1058-5591.

Scheifler:XR-10-1-13

[Sch94a] Bob Scheifler. From the X consortium: What’s new in R6. The X Resource, 10(1):13–31, April 1994. CODEN XRESEA. ISBN 1-56592-067-8. ISSN 1058-5591.

Scheifler:XR-12-1-15

[Sch94b] Bob Scheifler. From the X Consortium: X changes, changes X. The X Resource, 12(1):15–20, October 1994. CODEN XRESEA. ISBN 1-56592-069-4. ISSN 1058-5591.

Scheifler:XR-9-1-153

[Sch94c] Bob Scheifler. X Consortium, Inc.: A status report. The X Resource, 9(1):153, January 1994. CODEN XRESEA. ISBN 1-56592-066-X. ISSN 1058-5591.

Scheifler:XR-13-1-177

[Sch95a] Bob Scheifler. X Consortium status report. The X Resource, 13 (1):177, January 1995. CODEN XRESEA. ISBN 1-56592-121-6. ISSN 1058-5591.

Schmidgall:XR-14-1-59

[Sch95b] Jay D. Schmidgall. Tickling fvwm: Extending Tk as an fvwm module. The X Resource, 14(1): 59–87, April 1995. CODEN XRE-SEA. ISBN 1-56592-122-4. ISSN 1058-5591.

Schultz:XR-13-1-283

[Sch95c] Karl Schultz. What’s new in PEX and PEXlib 5.2. The X Resource, 13(1):283–292, January 1995. CODEN XRESEA. ISBN 1-56592-121-6. ISSN 1058-5591.

Shaffer:XR-6-1-79

[Sha93] David M. Shaffer. Interprocess communications in Xt programs. The X Resource, 6(1):79– 96, March 1993. CODEN XRE-SEA. ISBN 1-56592-021-X. ISSN 1058-5591.

Shaw:XR-9-1-151

[Sha94] Peter Shaw. The future of X. The X Resource, 9(1):151–152, January 1994. CODEN XRESEA. ISBN 1-56592-066-X. ISSN 1058-5591.

Solomita:XR-11-1-143

[SKD94] Ethan Solomita, James Kempf, and Dan Duchamp. XMOVE: A pseudoserver for X window movement. The X Resource, 11(1):143– 170, July 1994. CODEN XRE-SEA. ISBN 1-56592-068-6. ISSN 1058-5591.

Smith:XR-1-1-73

[Smi92] John Allen Smith. The multithreaded X server. The X Resource, 1(1):73–89, January 1992. CODEN XRESEA. ISBN 0-937175-96-X. ISSN 1058-5591.

Smyth:XR-1-2-158

[SN92] David E. Smyth and Adian Nye. Wcl 2.0: The widget creation library. The X Resource, 2(1):158– 182, April 1992. CODEN XRE-SEA. ISBN 0-937175-97-8. ISSN 1058-5591.

Srinivasan:XR-13-1-181

[Sri95] Murali V. Srinivasan. A methodology for multithreaded X client development. The X Resource, 13 (1):181, January 1995. CODEN XRESEA. ISBN 1-56592-121-6. ISSN 1058-5591.

Stabile:XR-15-1-43

[Sta95] Lawrence A. Stabile. Simplifying GUI construction by embedded scripting. The X Resource, 15(1):43–70, July 1995. CODEN XRESEA. ISBN 1-56592-140-2. ISSN 1058-5591.

Stevenson:XR-B-1-1b

[Ste92] Jeff Stevenson. PEXlib specification and C language binding. The X Resource, Special issue B(1):1– 464, September 1992. CODEN XRESEA. ISBN 1-56592-050-3. ISSN 1058-5591.

Stevenson:XR-6-1-115

[Ste93] Jeff Stevenson. PEX 6.0: Why?

What is it? And how will you get there? The X Resource, 6(1):115– 123, March 1993. CODEN XRE-SEA. ISBN 1-56592-021-X. ISSN 1058-5591.

Stevenson:XR-9-1-71

[Ste94] Jeff Stevenson. PEX 6.0 — NOT!

We’re doin’ 5.2 instead. The X Resource, 9(1):71, January 1994. CODEN XRESEA. ISBN 1-56592-066-X. ISSN 1058-5591.

Stevenson:XR-E-1-1

[Ste95] Jeff Stevenson. PEX protocol specification, version 5.2: Public review draft. The X Resource, Special issue E(1):xxiv + 603, October 1995. CODEN XRESEA. ISBN 1-56592-071-6. ISSN 1058-5591.

Strassl:XR-8-1-47

[Str93]  Bernhard Strassl.  Xm++—

another way to program the X toolkit. The X Resource, 8(1):47– 68, October 1993. CODEN XRE-SEA. ISBN 1-56592-023-6. ISSN 1058-5591.

Summers:XR-5-1-7

[SV93] Kent J. Summers and Jeffrey L. Vogel. Proposal for an X-based online help protocol. The X Resource, 5(1):7–24, January 1993. CODEN XRESEA. ISBN 1-56592-020-1. ISSN 1058-5591.

Smaalders:XR-1-1-90

[SWC92] Bart Smaalders, Brian Warkentine, and Kevin Clarke. Prototyping MT-safe Xt and XView libraries. The X Resource, 1(1): 90–107, January 1992. CODEN XRESEA. ISBN 0-937175-96-X. ISSN 1058-5591.

Szczur:XR-1-2-56

[Szc92] Marti Szczur. NASA’s TAE PLUS: a GUI development tool and application. The X Resource, 2(1):56–77, April 1992. CODEN XRESEA. ISBN 0-937175-97-8. ISSN 1058-5591.

Tabayoyon:XR-6-1-27

[Tab93] Al Tabayoyon. Resolving Xt resource collisions. The X Resource, 6(1):27–36, March 1993. CODEN XRESEA. ISBN 1-56592-021-X. ISSN 1058-5591.

Takano:XR-13-1-63

[TMM95] Hajime Takano, Hiroshi Matsuura, and Hiroshi Matoba. XVE: X visual effect extension. The X Resource, 13(1):63–74, January 1995. CODEN XRESEA. ISBN 1-56592-121-6. ISSN 1058-5591.

Tabayoyon:XR-0-1-61

[TT91] Al Tabayoyon and Joann M. Taylor. A technical introduction to the X Color Management System. The X Resource, 0(1):61–70, October 1991. CODEN XRESEA. ISBN 0-937175-79-X. ISSN 1058-5591.

VanHaren:XR-6-1-165

[Van93] Chris VanHaren. Dash: An application browser and help system. The X Resource, 6(1):165– 180, March 1993. CODEN XRE-SEA. ISBN 1-56592-021-X. ISSN 1058-5591.

Weinstein:XR-1-1-61

[Wei92] Jeff Weinstein. NFB, an X server porting layer. The X Resource, 1 (1):61–71, January 1992. CODEN XRESEA. ISBN 0-937175-96-X. ISSN 1058-5591.

Werner:XR-12-1-59

[Wer94] Rob Werner. The Xmsg library:

An application-level messagereporting facility. The X Resource, 12(1):59–91, October 1994. CODEN XRESEA. ISBN 1-56592-069-4. ISSN 1058-5591.

Wexler:XR-D-1-D

[Wex94a] Mike Wexler. X session management protocol: Public review draft: X consortium draft standard. The X Resource, Special issue D(1):D–i–D–iii, D–1–D–16, February 1994. CODEN XRE-SEA. ISBN 1-56592-071-6. ISSN 1058-5591.

Wexler:XR-9-1-163

[Wex94b] Mike Wexler. XSMP: The new session management protocol. The X Resource, 9(1):163– 166, January 1994. CODEN XRESEA. ISBN 1-56592-066-X. ISSN 1058-5591.

Womack:XR-A-1-1a

[WFB^{+}92] Paula Womack, Jeffrey Friedberg, Sally C. Barry, Randi J. Rost, and Todd Newman. PEX protocol encoding, version 5.1P—X public review draft. The X Resource, Special issue A(1):1–81, May 1992. CODEN XRESEA. ISBN 1-56592-024-4. ISSN 1058-5591.

Widener:XR-0-1-133

[WJ91] Glenn Widener and Vania Joloboff. Developing internationalized X clients: X11 Release 5 uses your language. The X Resource, 0(1): 133–153, October 1991. CODEN XRESEA. ISBN 0-937175-79-X. ISSN 1058-5591.

Walker:XR-5-1-213

[WNTV93] William D. Walker, Mark E. Novak, Henry R. Tumblin, and Gregg C. Vanderheiden. Making the X Window System accessible to people with disabilities. The X Resource, 5(1):213–227, January 1993. CODEN XRESEA. ISBN 1-56592-020-1. ISSN 1058-5591.

Womack:XR-A-1-1b

[Wom92] Paula Womack. PEX protocol specification and encoding, version 5.1P. The X Resource, Special issue A:1–179, May 1992. CO-DEN XRESEA. ISBN 1-56592-024-4. ISSN 1058-5591.

Yu:XR-9-1-237

[Yu94] Tom Yu. Kerberos authentication of X connections. The X Resource, 9(1):237–243, January 1994. CODEN XRESEA. ISBN 1-56592-066-X. ISSN 1058-5591.

Yoon:XR-15-1-85

[YYPC95] Jinsoo Yoon, Yanghee Yoon, Sunhgun Park, and Kilnam Chon. A Korean environment for the X Window System. The X Resource, 15(1):85–116, July 1995. CODEN XRESEA. ISBN 1-56592-140-2. ISSN 1058-5591.

Zimet:XR-9-1-229

[Zim94] Martha Zimet. Extending X for recording. The X Resource, 9(1): 229–235, January 1994. CODEN XRESEA. ISBN 1-56592-066-X. ISSN 1058-5591.

Zimet:XR-14-1-159

[Zim95a] Martha Zimet. Extending X for recording: Version 1.10 public review draft, 10 february 1995. The X Resource, 14(1):159–176, April 1995. CODEN XRESEA. ISBN 1-56592-122-4. ISSN 1058-5591.

Zimet:XR-14-1-177

[Zim95b] Martha Zimet. Record extension library specification: Version 1.10 public review draft, 10 february 1995. The X Resource, 14(1):177– 193, April 1995. CODEN XRE-SEA. ISBN 1-56592-122-4. ISSN 1058-5591.

Zimet:XR-16-1-103

[Zim95c] Martha Zimet. The X synchronization extension. The X Resource, 16(1):103–167, December 1995. CODEN XRESEA. ISBN 0-937175-79-X. ISSN 1058-5591.

Zimet:XR-15-1-119

[Zim95d] Martha Zimet. The X11 Testing Extension: A tutorial. The X Resource, 15(1):119–158, July 1995. CODEN XRESEA. ISBN 1-56592-140-2. ISSN 1058-5591.